DETERMINE RESISTANCEFROM CURRENT-VOLTAGE

Suppose you apply several different voltages across a circuit and measure the current that runs through the circuit. NOTE: the current is given in milliamperes. Example, at 2 voltage, the current is 0.002A or 2mA. What is the resistance of the circuit?


1
Expert's answer
2021-05-13T10:26:34-0400


We know that the relation is

"\boxed{V= IR}"

And

"\boxed{I =\frac{V}{R}}"

And

"\boxed{R=\frac{V}{I}}"

As given in example

V= 2

I =0.002 A

"R=\frac{V}{I}"

"R=\frac{2}{0.002}"

R(resistance )=1000 ohm